College of Computer & Information Science Spring 2004 Northeastern University Handout 1 CS G719: PhD Seminar 29 January 2004

Online Computation

Name:

1. Give a short summary of the seminar.

2. Describe one application scenario for online computation other than the ones discussed in the seminar. (Variants are ok.)

3. Ask a question.